If 1.00 g of element X reacts with 2.00 g of element Y to produce 3.00 g of a compound with the formula X2Y3, and the molar mass of Y is 50.0 g/mol, what must be the molar mass of X?

(a) First of all, after reading the question, which element do you know more about, X or Y?

(b) How many moles of Y are in 2.00 g of Y, and therefore must be in the X2Y3 ?

(c) Based on your answer to (b) above, and using the molar ratio within the compound X2Y3, how many moles of X must be in the X2Y3 ?

(d) Based on your answer to (c) above, what must be the molar mass of X ?  (keep in mind that to calculate a molar mass, you need two numbers: mass (g) and moles)
Answer #1

(a) Y. Because the molecular mass of Y is also mentioned.

(b) No: of moles = Mass (g) / molecular mass (g/mol)

So, No: of moles of Y (Ny) = 2/50 = 0.04

(c) In X2Y3, No: of moles of X (Nx) : No: of moles of Y (Ny) = 2:3

    Nx : 0.04 = 2:3

Nx / 0.04 = 2/3

Nx = 2* 0.04/3 = 0.08/3 = 0.027

(d) No: of moles = Mass (g) / molecular mass (g/mol)

So, molecular mass (g/mol) = Mass (g) / No: of moles

= 1/ (0.08/3)

= 3/0.08

= 37.5 g